OpenStudy (anonymous):

which expression gives the solution of -5+2x^2=-6x?

OpenStudy (anonymous):

\[2 +- \sqrt{4 - (4)(6)(-5)}\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[-5 +-\sqrt{25-(4)(2)(6)}\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[-6 +- \sqrt{36 -(4)(2)(-5)}\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[6 +- \sqrt{36 -(4)(2)(5)}/4\]

OpenStudy (anonymous):

the first one should be divided by 12 the second one should be divided by -10 the third one should be divided by 4

OpenStudy (kagıtucak):

the roots are from formulation  if you arrange the equation 2x^2+6x-5=0 here b=6 a=2 c=-5 and delta=b^2-4ac=36-4.2.(-5)=76 so roots are -6+76^1/2 over 4 and -6-76^1/2 over 4 so it the third one and like you said it should be divided by 4